Sol Summary – April 12th

Crew 178 Daily Summary Report 12 APR 2017
MDRS Daily Summary Report for sol 2
Summary Title: Second EVA, preparations for other experiments and Internet fixing
Mission Status: Good
Sol Activity Summary:
“Today, we again woke up at 8 am after having enjoyed “The Martian” movie yesterday night with the whole team. In the morning, Nathalie invented the “Pancake dance” while she was cooking delicious pancakes for everyone, although her first 5 pancakes were closer to sharcoal than cake. Elke finished the setup of the last parameters for her second EVA using her ground penetrating radar. Calogero continued to configure his software applications for his geomatics EVA. At noon, we ate strange canned turkey meat and a Minestrone bear creek soup with our favorite bread, oven-cooked with love by our Crew Scientist, Mathieu. On the whole, this meal was still a pleasure. In the afternoon, the second EVA was launched under a very sunny weather reaching about 26 degrees Celsius. A fun fact about our Biohazard space suits is that they’re cloaked with an external plastic layer, allowing us to sweat like never inside. Extreme conditions are fulfilled. Quentin and Elke will try to reconstruct a Shepherd’s pie with lyophilized food for an amazing dinner tonight. After dinner, every crew member will have still hard work to finish preparing at best the next days.
Aurian d’Avernas, the Crew Commander”
Look Ahead Plan: The HughesNet technician came today to fix the parable with concrete. We identified that some laptops were automatically burning our daily Internet data with the automatic Windows updates which are very tricky to disable. We finally found the solution and tomorrow will be the first day without any Internet issue. The preparations of the crew experiments are going well. Moreover, we will send our biographies and crew photos tomorrow morning without exception.
Anomalies in work: The wheel of the ground penetrating radar does not work properly. It should be fixed by a phone call to the company of this material tomorrow morning by our Crew Geologist.
Weather: Clear sky and very sunny.
Crew Physical Status: Very good.
EVA: 1 km at the North of the Hab.
Reports to be file: EVA #2 report + Journalist report + Engineer report + Daily Summary Report
Support Requested: none